OpenStudy (anonymous):

find all rational zeros of the polynomial: P(x) = x^4+x^3-x^2-2

OpenStudy (anonymous):

your options are \[\{\pm1,\pm2\}\]

OpenStudy (anonymous):

1 you can do by inspection \[1+1-1-2=-1\] no

OpenStudy (anonymous):

\[P(-1)=(-1)^4+(-1)^3-(-1)^2-2\] \[=1-1+1-2\] no

OpenStudy (anonymous):

maybe the other team can find one...

myininaya (myininaya):

no rational zeros

OpenStudy (anonymous):

beaten to the punch!

myininaya (myininaya):

satellite you are apart of our team

OpenStudy (anonymous):

\[\color{green}{\text{jealous}}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

<.<

myininaya (myininaya):

charles is joe's duck the secret is out

OpenStudy (anonymous):

ok 2 and - 2 don't work. no rational zeros

OpenStudy (anonymous):

lolol that sounds weird.

myininaya (myininaya):

lol

OpenStudy (anonymous):

Charles is swimming right now.

OpenStudy (anonymous):

they need to come up with a better problem than the fence built against the side of a river, barn, cliff, whatever.

OpenStudy (anonymous):

i think the river was a nice change of pace. you cant put a duck on a barn.

OpenStudy (anonymous):

or at least, a duck has no business on a barn.

myininaya (myininaya):

you do a pig

OpenStudy (anonymous):

im gonna start adding weir little things to all my drawings and see if people catch them.

myininaya (myininaya):

you can draw the pig eating charles